# Sorting inversion sequences

### v5

@@ Abstract
1.   Introduction
An inversion in a permutation $\pi=\pi_{1}\pi_{2}\pi_{3}\cdots\pi_{n}$ is a pair of entries $\pi_{i},\pi_{j}$ where $i<j$, but $\pi_{i}>\pi_{j}$. An inversion sequence of length $n$ is a word $e=e_{1}\dotsb e_{n}$ which satisfies for each $i\in\{1,\ldots,n\}$ the inequalities $10\leq e_{i}\leq i-$ 1. These sequences correspond to permutations via a variant of the Lehmer code [21, 22] by mapping each entry $\pi_{j}$ of the permutation $\pi$ to $e_{j-1}$ where $e_{j-1}$ is the number of inversions in $\pi$ where $\pi_{j}$ is the second, smaller entry of the pair. The set of inversion sequences (respectively the set of inversion sequences of length $n$) is denoted ${\mathbf{I}}$ (respectively ${\mathbf{I}}_{n}$).
@@ 1.   Introduction
The study of pattern avoidance in inversion sequences began with the work of Corteel, Martinez, Savage, and Weselcouch [12] as well as that of Shattuck and the first author [25]. Since then, extensive work has been done in this area, including [23, 24, 37] to list a few. Recently, Kotsireas, Yıldırım, and the first author [20] introduced an algorithmic technique involving generating trees to enumerate many pattern classes of inversion sequences (also see [7, 8] and the references therein). We utilize this algorithm in Section 4 to obtain a generating function for the number of inversion sequences sortable by two particular machines and give the succession rules to build the generating tree for a third machine.
@@
A stack is a last-in, first-out sorting device with push and pop operations. Knuth [19] showed the permutation $\pi$ can be sorted by a stack (that is, by applying push and pop operations to the sequence $\pi_{1},\dots,\pi_{n}$ one can output the identity permutation $1,\dots,n$) if and only if $\pi$ avoids the permutation $231$.
@@
A restricted version of a stack that has received a fair amount of attention is a pop stack, first introduced by Avis and Newborn [5]. A pop stack supports the same push and pop operations as a standard stack, but whenever a pop operation occurs, every entry in the pop stack is popped immediately. One foundational result is that the permutations that can be sorted by a single pop stack are exactly those that avoid $231$ and $312$. Some later works involving pop stacks include [1, 4, 9, 11, 13, 16, 18, 30, 34].
Atkinson [2] explored a variation of a stack called an $(r,s)$-stack which extends the traditional push and pop operations. Specifically, the first index of an $(r,s)$ stack allowed one to push an entry not only to the top of the stack but to any of the top $r$ spots. Similarly, the second index expanded the options of the pop operation to allow one to pop any of the top $s$ entries from the stack. Thus a $(1,1)$-stack is simply a traditional stack. Atkinson was able to characterize these permutations when one of $r,s$ was fixed at $1$, as well as in the case when $r=s=2$. Natural variations on stacks arise when studying inversion sequences that are similar in flavor to these generalized stacks.
Elder [14] considered restricting the depth of a stack, that is the maximum number of entries allowed in a stack at any given time, in the context of generating permutations with two stacks in series. Elder later continued this study with Lee and Rechnitzer [17] and then extended to sorting and pop stack sorting with Goh [15, 16]. Although this work focused on a permutation sorting (or generating) machine consisting of two stacks in series, we will consider a similar restriction for a single stack or pop stack when sorting inversion sequences.
@@
2.   Stack sorting inversion sequences
@@
The enumeration of these sortable inversion sequences has been studied, but thus far remains open. See [25, Section 4] and [12] for initial study and further work on avoiding $120$ in [37].
3.   Pop stack sorting inversion sequences
@@
We note that Cerbai [9, Theorem 5.2] found the basis for sortable Cayley words in the context of a hare pop stack which is simply a pop stack that only allows entries to appear in weakly decreasing order from top to bottom (which is a necessary requirement for sorting with a single stack). This basis is the same as the one in Theorem 3.1 below. Cerbai’s argument could also be extended to sorting general words by a pop stack as we have done here.
@@
At the other extreme, only one inversion sequence of any given length $n\geq 1$ can have $n$ layers, namely the increasing sequence $012\ldots(n-1)$. Reducing the number of layers by one to $n-1$ layers gives a set of inversion sequences enumerated by the Tetrahedral numbers, sequence number A000292 in OEIS [32].
@@
One last special case is when the sortable inversion sequence contains only 00s and $1$s and thus only needs to avoid $1010$, thereby having at most three layers. The number of such sequences is the number of compositions of $n$ into at most four parts because we must begin with a 00 and can only alternate values at most three times after that. These numbers (with an offset of one) are better known as the “Cake numbers" [36], sequence number A000125 in OEIS [32], which count the maximum number of slices of a cylindrical cake one can have after making $n-1$ planar cuts perpendicular to the top circular surface.
@@
We note that Cerbai [9] also considered a tortoise pop stack for sorting Cayley words that had an additional pattern restriction of not allowing a $00$ pattern to appear in the pop stack. (Restricting the content of stacks in terms of pattern avoidance was initially considered in [3, 33, 10].) Cerbai’s characterization of sortable Cayley words can be easily extended to all words, so we omit the proof here.
@@
Given this classification, the generating function for the enumeration of tortoise pop stack sortable inversion sequences is given in Theorem 12 in recent work by Callan and Mansour [8].
@@ 3.1 Restricting pop stack depth
When extending the notion of the depth of a stack to words, we could continue to use the definition in terms of the number of elements allowed in the stack as introduced by Elder [14]. This could be more practical in some applications. However, given the repetition that distinguishes words from permutations and considering the storage in terms of distinct elements to remember, it also makes sense to consider another definition that is still consistent with the original definition when applied to permutations. The content of the stack could be stored as a limited number (the depth) of lists of positions in the stack where these positions are counted from the bottom. We consider the latter notion as given in the following definition.
@@
In the case of depth one, the number of sortable inversion sequences is the number of weakly increasing inversion sequences which are known to be enumerated by the Catalan numbers due to Martinez and Savage [28]. The proof shows these inversion sequences to correspond to the $213$ avoiding permutations, which were enumerated by Simion and Schmidt [31].
@@
For a pop stack of depth two, the sortable inversion sequences would need to avoid $120,201,210,1010$. The first few terms of the enumeration sequence are: $1,1,2,6,23,100,471,2349$. We define the succession rules for the generating tree for these inversion sequences (using the algorithm introduced in [20]) in Section 4.3. However, it is hard to get an explicit formula by translating this generating tree to a system of equations and it is even more difficult for larger values of $k$.
@@
A weakly decreasing word of length $n$ has depth $k$ if it has the form $\{a_{1},...a_{1},a_{2},...a_{2},...,a_{k+1},...a_{k+1}\}$ where for each $i$, $a_{i}>a_{i+1}\geq 0$ (so there are $k$ descents). Let $WD_{k}(n,a)$ be the set of all such words whose first value is $a$ and whose depth is at most $k$.
@@
where $a_{1}=0$, and the $\sum^{*}$ means for all $j>1$ to only sum on values of $a_{j}$ such that ${0\leq a_{j}\leq\bigg{(}\sum_{h=1}^{j-1}m_{h}-a_{h}\bigg{)}}$.
@@
The depth 2 stack sortable words are enumerated by Burstein [6] in the context of avoiding $120,210$.
@@
Finally, we note that Kotsireas, Yıldırım, and the first author [20] give a functional equation for the inversion sequences of length $n$ avoiding $120,210$ and thereby sortable by a stack of depth $2$.
@@ 3.2 Generalizing pop stacks
We can also expand the ability of our sorting machine using a similarly defined version of Atkinson’s [2] extension of push and pop operations. Specifically, we will define a $(r,1)$-pop stack that is also modified to take into account the repetition of letters in inversion sequences (and more generally words). In Atkinson’s work, an $(r,1)$-stack would have allowed for push operations to push entries into the stack in any of the top $r$ positions. We can again naturally redefine generalized stacks for words in a way that is consistent with Atkinson’s work on permutations.
@@
The enumeration problem for this $(2,1)$-pop stack of depth $2$ can be considered for permutations, words, and specifically inversion sequences. The permutations of length $n$ avoiding $231,312,321$, i.e. $120,201,210$, are known as the layered permutations where each layer can have size one or two. Note that layered permutations are exactly the permutations that avoid $231,312$. Furthermore, the restriction of avoiding $321$ forces each layer (consecutive decreasing subsequence) to have size at most two. These restricted layered permutations are enumerated by the Fibonacci numbers as shown by Simion and Schmidt [31]. As a consequence of their result and Theorem 3.23 we have the following corollary.
@@
Burstein [6, Theorem 5.1] enumerated the words of length $n$ over an alphabet $[k]$ avoiding $120,201,210$.
@@
4.   The generating tree method
@@
To enumerate the inversion sequences of length $n$ avoiding $120,201,210$, i.e. the inversion sequences sortable by a $(2,1)$-pop stack of depth $2$, we use the generating tree method shown by Kotsireas, Yıldırım, and the first author [20]. See the recent paper of Pantone [29] for another nice description of using succession rules to enumerate pattern avoiding inversion sequences.
@@
A succession rule, denoted $x\rightsquigarrow y_{1},y_{2},\ldots,y_{k}$, will indicate that the set $\{y_{1},y_{2},\ldots,y_{k}\}$ will be such that the generating subtrees rooted at $y_{1},y_{2},\ldots,y_{k}$ are equivalent to those rooted at $z_{1},z_{2},\ldots,z_{k}$ where $\{z_{1},z_{2},\ldots,z_{k}\}$ is the complete set of children of $x$.
@@
Also, the children of $b_{m,j}\in{\mathbf{T}}^{\prime}$ are $b_{m,j}0,b_{m,j}1,\ldots,b_{m,j}(m+1)$. Notice that ${\mathbf{T}}(b_{m,j}k)\cong{\mathbf{T}}(b_{m+2-j,1})$ with $k=0,1,\ldots,j-1$. To show this, we map any inversion sequence $\pi=0^{m}jk\pi^{\prime}\in{\mathbf{I}}_{n}(120,201,210)$ to $0^{m+2-j}1\pi^{\prime\prime}$, where $\pi^{\prime\prime}$ is obtained from $\pi^{\prime}$ (there are no letters in $\pi^{\prime}$ belonging to the set $\{0,\ldots,k-1,k+1,\ldots,j-1\}$) by replacing each letter $x\geq j$ by $x+1-j$ and replacing the letter $k$ by 00. Hence, we see that $\pi\in{\mathbf{I}}_{n}(120,201,210)$ if and only if $0^{m+2-j}1\pi^{\prime\prime}\in I_{n+2-j}(120,201,210)$.
@@ 4.2 The generating tree for the pop stack sortable inversion sequences
As in the previous case, one can show that the generating tree (based on the algorithm given in [20]) of $I_{n}(120,201,1010)$, that is, of the pop stack sortable inversion sequences, is given by a root $a_{1}$ and the rules
@@ 4.3 Inversion sequences sortable by a pop stack of depth two
where $a_{m}=0^{m},b_{m,j}=0^{m}j$, and $c_{m}=0^{m}10$. We label the inversion sequence 00 by $a_{1}$, so the root of the tree $T_{P}$ is indeed $a_{1}$. Now, let us show that the succession rules hold in $T_{P}$. Based on the algorithm introduced in [17] (for more examples and proofs, see [7, 8]), we have the following:
@@
5.   Open Problems
There are many directions left to continue the study of sortable inversion sequences. The most obvious question is the elusive enumeration problem for the inversion sequences of length $n$ which avoid $120$. Already interesting as a standalone pattern avoidance question, the fact that these inversion sequences are exactly those sortable by the classic stack sorting algorithm by Knuth [19] increases the intrigue. The first author and Shattuck [25, Section 4] obtained a recurrence relation for these inversion sequences yielding a functional equation in three variables (sequence number A263778 in OEIS [32]). We note that the construction method used here was later generalized by Testart [35] as part of his work to complete the enumeration of 22 open cases of pattern-avoiding inversion sequences.

### v6

@@ Abstract
keywords:
inversion sequence, permutation pattern, sorting algorithm, pop stack, generating tree
1 Introduction
An inversion in a permutation $\pi=\pi_{1}\pi_{2}\pi_{3}\cdots\pi_{n}$ is a pair of entries $\pi_{i},\pi_{j}$ where $i<j$, but $\pi_{i}>\pi_{j}$. An inversion sequence of length $n$ is a word $e=e_{1}\ldots e_{n}$ which satisfies for each $i\in\{1,\ldots,n\}$ the inequalities $10\leq e_{i}\leq i-$ 1. These sequences correspond to permutations via a variant of the Lehmer code Laisant (1888); Lehmer (1960) by mapping each entry $\pi_{j}$ of the permutation $\pi$ to $e_{j-1}$ where $e_{j-1}$ is the number of inversions in $\pi$ where $\pi_{j}$ is the second, smaller entry of the pair. The set of inversion sequences (respectively the set of inversion sequences of length $n$) is denoted ${\mathbf{I}}$ (respectively ${\mathbf{I}}_{n}$).
@@ 1 Introduction
The study of pattern avoidance in inversion sequences began with the work of Corteel, Martinez, Savage, and Weselcouch 2016 as well as that of Shattuck and the first author 2015. Since then, extensive work has been done in this area, including Lin (2020); Lin and Kim (2021); Yan and Lin (2020–2021) to list a few. Recently, Kotsireas, Yıldırım, and the first author 2024 introduced an algorithmic technique involving generating trees to enumerate many pattern classes of inversion sequences (also see Callan et al. (2023); Callan and Mansour (2023) and the references therein). We utilize this algorithm in Section 4 to obtain a generating function for the number of inversion sequences sortable by two particular machines and give the succession rules to build the generating tree for a third machine.
@@
A stack is a last-in, first-out sorting device with push and pop operations. Knuth 1969 showed the permutation $\pi$ can be sorted by a stack (that is, by applying push and pop operations to the sequence $\pi_{1},\dots,\pi_{n}$ one can output the identity permutation $1,\dots,n$) if and only if $\pi$ avoids the permutation $231$.
@@
A restricted version of a stack that has received a fair amount of attention is a pop stack, first introduced by Avis and Newborn 1981. A pop stack supports the same push and pop operations as a standard stack, but whenever a pop operation occurs, every entry in the pop stack is popped immediately. One foundational result is that the permutations that can be sorted by a single pop stack are exactly those that avoid $231$ and $312$. Some later works involving pop stacks include Asinowski et al. (2019); Atkinson and Sack (1999); Cerbai (2021); Claesson and Guðmundsson (2019); Defant and Williams (2022); Elder and Goh (2021); Hong (2022); Pudwell and Smith (2019); Smith and Vatter (2009).
Atkinson 1998 explored a variation of a stack called an $(r,s)$-stack which extends the traditional push and pop operations. Specifically, the first index of an $(r,s)$ stack allowed one to push an entry not only to the top of the stack but to any of the top $r$ spots. Similarly, the second index expanded the options of the pop operation to allow one to pop any of the top $s$ entries from the stack. Thus a $(1,1)$-stack is simply a traditional stack. Atkinson was able to characterize these permutations when one of $r,s$ was fixed at $1$, as well as in the case when $r=s=2$. Natural variations on stacks arise when studying inversion sequences that are similar in flavor to these generalized stacks.
Elder 2006 considered restricting the depth of a stack, that is the maximum number of entries allowed in a stack at any given time, in the context of generating permutations with two stacks in series. Elder later continued this study with Lee and Rechnitzer 2015 and then extended to sorting and pop stack sorting with Goh 2018; 2021. Although this work focused on a permutation sorting (or generating) machine consisting of two stacks in series, we will consider a similar restriction for a single stack or pop stack when sorting inversion sequences.
@@
2 Stack sorting inversion sequences
@@
The enumeration of these sortable inversion sequences has been studied, but thus far remains open. See (Mansour and Shattuck, 2015, Section 4) and Corteel et al. (2016) for initial study and further work on avoiding $120$ in Yan and Lin (2020–2021).
3 Pop stack sorting inversion sequences
@@
We note that Cerbai 2021, Theorem 5.2 found the basis for sortable Cayley words in the context of a hare pop stack which is simply a pop stack that only allows entries to appear in weakly decreasing order from top to bottom (which is a necessary requirement for sorting with a single stack). This basis is the same as the one in Theorem 3.1 below. Cerbai’s argument could also be extended to sorting general words by a pop stack as we have done here.
@@
At the other extreme, only one inversion sequence of any given length $n\geq 1$ can have $n$ layers, namely the increasing sequence $012\ldots(n-1)$. Reducing the number of layers by one to $n-1$ layers gives a set of inversion sequences enumerated by the Tetrahedral numbers, sequence number A000292 in OEIS Sloane (2025).
@@
One last special case is when the sortable inversion sequence contains only 00s and $1$s and thus only needs to avoid $1010$, thereby having at most three layers. The number of such sequences is the number of compositions of $n$ into at most four parts because we must begin with a 00 and can only alternate values at most three times after that. These numbers (with an offset of one) are better known as the “Cake numbers" Yaglom and Yaglom (1987), sequence number A000125 in OEIS Sloane (2025), which count the maximum number of slices of a cylindrical cake one can have after making $n-1$ planar cuts perpendicular to the top circular surface.
@@
We note that Cerbai 2021 also considered a tortoise pop stack for sorting Cayley words that had an additional pattern restriction of not allowing a $00$ pattern to appear in the pop stack. (Restricting the content of stacks in terms of pattern avoidance was initially considered in Atkinson et al. (2002); Smith (2014); Cerbai et al. (2020).) Cerbai’s characterization of sortable Cayley words can be easily extended to all words, so we omit the proof here.
@@
Given this classification, the generating function for the enumeration of tortoise pop stack sortable inversion sequences is given in Theorem 12 in recent work by Callan and Mansour 2023.
@@ 3.1 Restricting pop stack depth
When extending the notion of the depth of a stack to words, we could continue to use the definition in terms of the number of elements allowed in the stack as introduced by Elder 2006. This could be more practical in some applications. However, given the repetition that distinguishes words from permutations and considering the storage in terms of distinct elements to remember, it also makes sense to consider another definition that is still consistent with the original definition when applied to permutations. The content of the stack could be stored as a limited number (the depth) of lists of positions in the stack where these positions are counted from the bottom. We consider the latter notion as given in the following definition.
@@
In the case of depth one, the number of sortable inversion sequences is the number of weakly increasing inversion sequences which are known to be enumerated by the Catalan numbers due to Martinez and Savage 2018. The proof shows these inversion sequences to correspond to the $213$ avoiding permutations, which were enumerated by Simion and Schmidt 1985.
@@
For a pop stack of depth two, the sortable inversion sequences would need to avoid $120,201,210,1010$. The first few terms of the enumeration sequence are: $1,1,2,6,23,100,471,2349$. We define the succession rules for the generating tree for these inversion sequences (using the algorithm introduced in Kotsireas et al. (2024)) in Section 4.3. However, it is hard to get an explicit formula by translating this generating tree to a system of equations and it is even more difficult for larger values of $k$.
@@
A weakly decreasing word of length $n$ has depth $k$ if it has the form $a_{1}\ldots a_{1}a_{2}\ldots a_{2}\ldots a_{k+1}\ldots a_{k+1}$ where for each $i$, $a_{i}>a_{i+1}\geq 0$ (so there are $k$ descents). Let $WD_{k}(n,a)$ be the set of all such words whose first value is $a$ and whose depth is at most $k$.
@@
where $a_{1}=0$, and $\sum^{*}$ means for all $j>1$ to sum on values of $a_{j}$ such that ${0\leq a_{j}\leq\bigg{(}\sum_{h=1}^{j-1}m_{h}-a_{h}\bigg{)}}$.
@@
The depth 2 stack sortable words are enumerated by Burstein 1988 in the context of avoiding $120,210$.
@@
Finally, we note that Kotsireas, Yıldırım, and the first author 2024 give a functional equation for the inversion sequences of length $n$ avoiding $120,210$ and thereby sortable by a stack of depth $2$.
@@ 3.2 Generalizing pop stacks
We can also expand the ability of our sorting machine using a similarly defined version of Atkinson’s 1998 extension of push and pop operations. Specifically, we will define a $(r,1)$-pop stack that is also modified to take into account the repetition of letters in inversion sequences (and more generally words). In Atkinson’s work, an $(r,1)$-stack would have allowed for push operations to push entries into the stack in any of the top $r$ positions. We can again naturally redefine generalized stacks for words in a way that is consistent with Atkinson’s work on permutations.
@@
The enumeration problem for this $(2,1)$-pop stack of depth $2$ can be considered for permutations, words, and specifically inversion sequences. The permutations of length $n$ avoiding $231,312,321$, i.e. $120,201,210$, are known as the layered permutations where each layer can have size one or two. Note that layered permutations are exactly the permutations that avoid $231,312$. Furthermore, the restriction of avoiding $321$ forces each layer (consecutive decreasing subsequence) to have size at most two. These restricted layered permutations are enumerated by the Fibonacci numbers as shown by Simion and Schmidt 1985. As a consequence of their result and Theorem 3.23 we have the following corollary.
@@
Burstein 1988, Theorem 5.1 enumerated the words of length $n$ over an alphabet $[k]$ avoiding $120,201,210$.
@@
4 The generating tree method
@@
To enumerate the inversion sequences of length $n$ avoiding $120,201,210$, i.e. the inversion sequences sortable by a $(2,1)$-pop stack of depth $2$, we use the generating tree method shown by Kotsireas, Yıldırım, and the first author 2024. See the recent paper of Pantone 2024 for another nice description of using succession rules to enumerate pattern avoiding inversion sequences.
@@
A succession rule, denoted $x\rightsquigarrow y_{1},y_{2},\ldots,y_{k}$, will indicate that the set $\{y_{1},y_{2},\ldots,y_{k}\}$ will be such that the generating subtrees rooted at $y_{1},y_{2},\ldots,y_{k}$ are equivalent to those rooted at
$z_{1},z_{2},\ldots,z_{k}$ where $\{z_{1},z_{2},\ldots,z_{k}\}$ is the complete set of children of $x$.
@@
Also, the children of $b_{m,j}\in{\mathbf{T}}^{\prime}$ are $b_{m,j}0,b_{m,j}1,\ldots,b_{m,j}(m+1)$. Notice ${\mathbf{T}}(b_{m,j}k)\cong{\mathbf{T}}(b_{m+2-j,1})$ with $k=0,1,\ldots,j-1$. To show this, we map any inversion sequence $\pi=0^{m}jk\pi^{\prime}\in{\mathbf{I}}_{n}(120,201,210)$ to $0^{m+2-j}1\pi^{\prime\prime}$, where $\pi^{\prime\prime}$ is obtained from $\pi^{\prime}$ (there are no letters in $\pi^{\prime}$ belonging to the set $\{0,\ldots,k-1,k+1,\ldots,j-1\}$) by replacing each letter $x\geq j$ by $x+1-j$ and replacing the letter $k$ by 00. Hence, we see that $\pi\in{\mathbf{I}}_{n}(120,201,210)$ if and only if $0^{m+2-j}1\pi^{\prime\prime}\in I_{n+2-j}(120,201,210)$.
@@ 4.2 The generating tree for the pop stack sortable inversion sequences
As in the previous case, one can show that the generating tree (based on the algorithm given in Kotsireas et al. (2024)) of $I_{n}(120,201,1010)$, that is, of the pop stack sortable inversion sequences, is given by a root $a_{1}$ and the rules
@@ 4.3 Inversion sequences sortable by a pop stack of depth two
where $a_{m}=0^{m},b_{m,j}=0^{m}j$, and $c_{m}=0^{m}10$. We label the inversion sequence 00 by $a_{1}$, so the root of the tree $T_{P}$ is indeed $a_{1}$. Now, let us show that the succession rules hold in $T_{P}$. Based on the algorithm introduced in [17] (for more examples and proofs, see Callan et al. (2023); Callan and Mansour (2023)), we have the following:
@@
5 Open Problems
There are many directions left to continue the study of sortable inversion sequences. The most obvious question is the elusive enumeration problem for the inversion sequences of length $n$ which avoid $120$. Already interesting as a standalone pattern avoidance question, the fact that these inversion sequences are exactly those sortable by the classic stack sorting algorithm by Knuth 1969 increases the intrigue. The first author and Shattuck 2015, Section 4 obtained a recurrence relation for these inversion sequences yielding a functional equation in three variables (sequence number A263778 in OEIS Sloane (2025)). We note that the construction method used here was later generalized by Testart 2024 as part of his work to complete the enumeration of 22 open cases of pattern-avoiding inversion sequences.